how many moles of sodium oxide are produced if 0.5 mol of Fe is produced in the reaction 6Na + Fe2O3 3Na2O + 2Fe


1
Expert's answer
2021-04-21T03:53:58-0400

6Na + Fe2O3 = 3Na2O + 2Fe

n (Na2O) = 3/2 n (Fe) = 3/2 x 0.5 = 0.33 mol